hallm:

solve the equation for x 9/2 (8-x) +36 =102 - 5/2 (3x +24) please dont just solve it for me explain it too

Karla:

In these types of questions you have to separate x /take x on one side of equation to know what is is value okay So first in order to remove denominator 2 multiply both right hand side as well as left hand side with 2 As shown below 2*( 9/2(8-x)+36) = 2*(102-5/2(3x+24) So it's same as 2*9/2*(8-x) + 36*2 =2*102 - 2*(5/2(3x+24)) Solve it 9(8-x)+72=204- 5(3x+24) Multiply 72-9x+72= 204 -(15x+120) For right hand side that is 204 -(15x+120) When you will open the bracket as product of negative and positive is negative It will be 204 -15x-120 on right hand side Coming back to whole equation 72+72-9x=204-120-15x 144-9x=84-15x Take x on one side and number on other this is done so that you can separate x and find its value So In order to take 15x on other side add 15x on both sides 144-9x+15x =84-15x+15x On right hand side -15x+15x=0 So it will be 144+6x=84 Take 144 in other side by subtracting it from both sides 144-144+6x=84-144 On left hand side 144-144=0 So it is 6x=-60 In order to separate x Divide both side by 6 6/6x = -60/6 On left hand side 6/6=1 So x=-10 So your answer for x is -10 Hope that helps

Karla:

https://youtu.be/p5e5mf_G3FI may be this will help

jhonyy9:

@Karla this is a very very nice job --- CONGRATS !!!

Karla:

Thank you :)